Find aV6 for aVn=8(1/2)^n, n=1,2,3,4,... Please help me to solve this problem.
Answer by venugopalramana(3286)

Find aV6 for aVn=8(1/2)^n, n=1,2,3,4,... Please help me to solve this problem.
IT APPEARS A=8 AND VN=(1/2)^N...IF SO WE HAVE FOR
N=1.......AV1=8(1/2)=4
N=2.......AV2=8(1/2)^2=8/4=2
N=3.......AV3=8(1/2)^3=8/8=1
N=4.......AV4=8(1/2)^4=8/16=1/2.....ETC